# Characterizing graph-nonedge pairs with single interval Cayley configuration spaces in 3-dimensions

### v1

@@
Abstract.
A linkage $(G,\ell)$ is a pair containing a finite simple undirected graph $G$ and a squared edge-length map $\ell$ that assigns squared Euclidean lengths to the edges of $G$. A $d$-realization of $(G,\ell)$ is an assignment of points in $\mathbb{R}^{d}$ to the vertices of $G$ for which pair-wise distances between points agree with $\ell$. For $d\leq 3$, we characterize pairs $(G,f)$, where $f$ is a nonedge of $G$, such that, for any squared edge-length map $\ell$, there is a single interval of attained distance values between the endpoints of $f$ over all $d$-realizations of $(G,\ell)$, answering a question posed in [22] a decade ago, which gave an equivalent characterization for $d\leq 2$ that does not generalize to $d\geq 3$. Two notable byproducts of this characterization are a new tool for partial $3$-tree completion, a well-studied problem, and a deeper understanding of the realization space of partial $3$-tree linkages through the lens of Cayley realization spaces. Although related to the minor closed class of $3$-flattenable graphs, the class of pairs $(G,f)$ with the above property is not minor closed, has no obvious well quasi-ordering, and there are infinitely many minimal graphs-nonedge pairs - w.r.t. edge contractions - in the complement class. Our characterization overcomes these obstacles, is based on the forbidden minors of the class of $3$-flattenable graphs, and contributes to the theory of Cayley configurations used for analyzing distance-constrained configuration spaces in a range of application domains. Generalizations to higher dimensions and efficient algorithmic characterizations are conjectured.
Key words and phrases:
@@ Abstract.
Configuration spaces, Graph flattenability, Cayley configuration spaces, Forbidden minor characterization, Distance geometry, Convex geometry
1CISE Department, University of Florida, partially supported by NSF DMS 1564480, NSF DMS 1563234, and DARPA HR00111720031.
2Affiliate in Department of Mathematics
1. Introduction
@@
In this paper, all sets are closed and compact, all maps are closed, and connected implies path-connected unless otherwise specified. A linkage $(G,\ell)$ is a pair containing a (finite simple undirected) graph $G$ and a squared edge-length map $\ell$ that sends each edge in the edge set $E(G)$ to a nonnegative real number. For any integer $d\geq 0$, a (Euclidean) $d$-realization $p$ of $(G,\ell)$ is a map that sends each vertex in the vertex set $V(G)$ to a point in $\mathbb{R}^{d}$ such that $\ell(uv)=\|p(u)-p(v)\|^{2}$ for each edge $uv\in E(G)$. We refer to $p$ as a realization of $(G,\ell)$ when $d$ is clear from context. See Figures 1 and 2 for examples. The configuration space (CS) $\mathcal{C}^{d}(G,\ell)$ is the set of all $d$-realizations of $(G,\ell)$. Traversing and sampling $\mathcal{C}^{d}(G,\ell)$ are difficult and well-studied problems [10, 11, 14, 21, 23, 26, 30] with applications in many domains [16, 18, 29].
Figure 1. (A) A linkage $(G,\ell)$, where $u_{1}u_{2}$ is a nonedge, and (B) a $2$-realization of $(G,\ell)$, where $p_{i}=p(v_{i})$. (A), (C), and (D) show graph-nonedge pairs. See the discussion below.
@@
Figure 2. Left: linkages $(G,\ell)$ and sets $F$ of nonedges (dashed line-segments). Top-right: the CCS $\Omega^{3}_{F}(G,\ell)$ for the top-left linkage; credit to [27]. Bottom-right: the CCS $\Omega^{2}_{F}(G,\ell)$ for the bottom-left linkage. Credit for both bottom figures to [22].
@@
The following concepts were introduced by Sitharam and Gao [22] to study these problems. Letting $F=\{u_{1}v_{1},\dots,u_{m}v_{m}\}$ be an indexed set of nonedges - i.e., pairs of distinct vertices not in $E(G)$ of $G$, the Cayley map $\phi_{F}$ sends a realization $p$ in $\mathcal{C}^{d}(G,\ell)$ to the vector $(\|p(u_{1})-p(v_{1})\|^{2},\dots,\|p(u_{m})-p(v_{m})\|^{2})$ of squared lengths attained by $F$ under $p$. The Cayley configuration space (CCS) $\Omega^{d}_{F}(G,\ell)$ is the image of $\mathcal{C}^{d}(G,\ell)$ under $\phi_{F}$. Note that the preimage of any point in $\Omega^{d}_{F}(G,\ell)$ under $\phi_{F}$ is a non-empty subset of $\mathcal{C}^{d}(G,\ell)$. See Figure 2 for examples. We call $(G,F)$ a graph-nonedge-set pair. If $F$ contains exactly one nonedge $f$, then we call $(G,f)$ a graph-nonedge pair. The pair $(G,F)$ is $d$-convex (resp. $d$-Cayley-connected) if $\Omega^{d}_{F}(G,\ell)$ is convex (resp. connected) for any squared edge-length map $\ell$. Observe that $(G,F)$ is $d$-convex and $d$-Cayley-connected if $d=0$. Hence, we assume that $d\geq 1$ unless stated otherwise. We study the following property.
@@
Given an integer $d\geq 0$, a graph-nonedge pair $(G,f)$ has the $d$-single interval property (SIP) if it is $d$-convex (equivalently, $d$-Cayley-connected).
1.0.1. Example
Refer to the graph-nonedge pairs $(G,f)$ in Figure 1. For the linkage $(G,\ell)$ in Figure 1(a), observe that the CS $\mathcal{C}^{1}(G,\ell)$ contains exactly two realizations up to translation and overall reflection, and the CCS $\Omega^{1}_{f}(G,\ell)$ is the set $\{1,9\}$. Hence, $(G,f)$ does not have the $1$-SIP. On the other hand, for any $\ell$ and integer $d\geq 2$, the triangle-inequality shows that the CCS $\Omega^{d}_{f}(G,\ell)$ is a single interval, and hence $(G,f)$ has the $d$-SIP. The $2$-realization in Figure 1(b) provides intuition for why $\Omega^{2}_{f}(G,\ell)$ is a single interval. For similar reasons, in Figure 1(c), $(G,f)$ has the $d$-SIP for any $d\geq 3$ but not for any $d<3$ and, in Figure 1(d), $(G,f)$ has the $d$-SIP for any $d\geq 4$ but not for any $d<4$.
@@ 1.0.1. Example
Next, refer to the graph-nonedge-sets $(G,F)$ in Figure 2. For the linkage $(G,\ell)$ in the bottom-left image, the triangle-inequality can be used to compute the boundaries of the CCS $\Omega^{2}_{F}(G,\ell)$ shown in the bottom-right image, which is convex. The convexity of this CCS depends only on the fact that it is the intersection of half-spaces, and not on the precise map $\ell$, and so $(G,F)$ is $d$-convex for any $d\geq 2$. By choosing $\ell$ so that it assigns one edge 00 and the rest $1$, it is easy to see that $\Omega^{1}_{f}(G,\ell)$ has cardinality two, and so $(G,F)$ is not $1$-convex. In the top-left image, similar argument using the tetrahedral-inequality shows that $(G,F)$ is $d$-convex for any $d\geq 3$ but not for any $d<3$. The top-right image shows the convex CCS $\Omega^{3}_{F}(G,\ell)$ for the linkage $(G,\ell)$ in the top-left image. It is easy to see that the CS $\mathcal{C}^{2}(G,\ell)$ is connected for any $\ell$, and so $\Omega^{3}_{F}(G,\ell)$ is connected for any $\ell$ since $\phi_{F}$ is continuous. Therefore, $(G,F)$ is $2$-Cayley-connected. However, an argument similar to those above shows that $(G,F)$ is not $1$-Cayley-connected.
@@
Figure 3. Three $d$-convex graph-nonedge-set pairs (nonedges are shown as dashed line-segments) for any $d\geq 3$. The left and middle graphs together with their non-edge-sets are the graphs $C_{5}\times C_{2}$ and $V_{8}$, respectively. See Examples 1.0.1 and 2.0.1 and Section 1.2.
@@
Finally, the graph-nonedge-set pairs $(G,F)$ in Figure 3 are $d$-convex for any $d\geq 3$ but not for any $d<3$. The former fact can be verified for the first two pairs using Theorem 3 in Section 1.2 and for the third pair using Theorem 4 in Section 2 - our main result. The latter fact can be verified using Theorems 1 and 2 in Section 1.2.
1.1. Organization
@@
Previous results on $d$-conxevity, $d$-Cayley-connectedness, and the $d$-SIP are collected in Section 1.2. We state our main results in Section 2. The majority of this paper is spent proving Theorem 4, our main result, which gives a characterization of graph-nonedge pairs that have the $3$-SIP. The primary proof obstacles are detailed in Section 2.1. The proof is given in Sections 3 (converse direction) and 4 (forward direction). Finally, we state several open problems in Section 5, including a conjectured characterization of graph-nonedge pairs that have the $d$-SIP for any $d\geq 1$.
1.2. Related Works
Theorems 1 and 2, below, are characterizations of graph-nonedge-set pairs that are $d$-convex and $d$-SIP for any $d\leq 2$, given by Sitharam and Gao [22] and restated using our terminology. The proof of the former strongly relies on the latter. See Example 1.0.1, in Section 1, and Example 1.2.1, below, for example applications of these theorems. We require the following concepts to state these theorems.
Figure 4. (A) a graph-nonedge pair $(G,f)$ with a clique separator $C$, shown in purple. (B) the $C$-components of $G\cup f$. (C) the atoms of $G\cup f$. See below, Example 2.0.1 in Section 2, and Example 4.0.1 in Section 4.
@@ 1.2. Related Works
Given a graph $G$ and a set of its edges or nonedges $F$, $G\cup F$ is the graph obtained from $G$ by adding $F$ to $E(G)$. For any subset $U\subseteq V(G)$, $G\setminus U$ is the graph obtained from $G$ by removing $U$ from $V(G)$. If either $F$ or $U$ contains exactly one element $f$ or $u$, respectively, then we replace $F$ with $f$ and $U$ with $u$ in this notation. A subgraph $H$ of $G$ is induced if, for any $u,v\in V(H)$, $uv\in E(H)$ if and only if $uv\in E(G)$. We say that $H$ is induced by $V(H)$. $U$ is a separator of $G$ if $G\setminus U$ is disconnected, a clique of $G$ if its induces a clique subgraph, and a clique separator (cs) of $G$ if it is both a clique and a separator. An atom is a graph with no cs. An atom of $G$ is a vertex-maximal induced subgraph that has no cs. See Figure 4 for examples. A graph $M$ is a minor of $G$ if it is isomorphic to some graph that can be obtained from $G$ via some sequence of vertex deletions, edge deletions, and edge contractions - i.e., deleting two vertices connected by an edge and adding a new vertex whose neighborhood is the union of the neighborhoods of the deleted vertices.
@@
For any integer $d\leq 2$, a graph-nonedge-set pair $(G,F)$ is $d$-convex if and only if no atom of $G\cup F$ that contains any non-empty subset of $F$ has a $K_{d+2}$ minor.
@@
For any integer $d\leq 2$, a graph-nonedge pair $(G,f)$ has the $d$-SIP if and only if no atom of $G\cup f$ that contains $f$ has a $K_{d+2}$ minor.
Figure 5. The forbidden minors for $3$-flattenability: $K_{5}$ (left) and $K_{2,2,2}$ (right). See below.
There are several known connections between $d$-convexity and a property of graphs called $d$-flattenability, which was introduced by Belk and Connelly [5, 4]. A graph $G$ is $d$-flattenable if, for any squared edge-length map $\ell$ and any integer $n\geq d$, the CS $\mathcal{C}^{d}(G,\ell)$ is non-empty whenever the CS $\mathcal{C}^{n}(G,\ell)$ is non-empty. Intuitively, this means that any high-dimensional realization of any linkage $(G,\ell)$ can be flattened to live in $d$-dimensions while preserving squared edge-lengths. For example, $K_{3}$ is $2$-flattenable but not $1$-flattenable, since choosing a squared edge-length map $\ell$ that assigns $1$ to each edge of $K_{3}$ ensures the CS $\mathcal{C}^{1}(K_{3},\ell)$ is empty while the CS $\mathcal{C}^{2}(K_{3},\ell)$ is not.
A property of graphs is minor closed if the family of graphs that have this property is closed under edge deletions and contractions. It is easy to see that $d$-flattenability is minor closed. Alternatively, if a graph is $d$-flattenable, then so are all of its minors. Hence, Robertson and Seymour’s famous Graph Minor Theorem [17] states that there exists a finite set of forbidden minors such that a graph is $d$-flattenable if and only if it has no minor in this set. Belk and Connelly [5, 4] showed that this set contains only $K_{3}$ for $d=2$, $K_{4}$ for $d=3$, and $K_{5}$ and $K_{2,2,2}$ for $d=3$ (see Figure 5). Finite forbidden minor characterizations for $d$-flattenability have also been studied for normed spaces other than Euclidean [20, 9, 8].
With these characterizations, the right-hand side of Theorem 2 can be replaced with “no atom of $G\cup f$ that contains $f$ has a $d$-flattenability forbidden minor”. We will see in Example 2.0.1, in Section 2, that this restatement does not extend to $d=3$. Sitharam and Gao [22] established the following relationship between $d$-convexity and $d$-flattenability for $d=3$, and Sitharam and Willoughby [25] later extended it to any $d\geq 0$ and to other norms. This relationship makes use of previous work on understanding the Euclidean and non-Euclidean distance cones [3, 19]. Given a graph $G$ and a set $F$ of its edges and nonedges, $G\setminus F$ is the graph obtained from $G$ by removing $F$ from $E(G)$. If $F$ contains exactly one element $f$, then we replace $F$ with $f$ in this notation.
@@
Consider a graph $G$ and an integer $d\geq 0$. For any subset $F\subseteq E(G)$, the pair $(G\setminus F,F)$ is $d$-convex if and only if $G$ is $d$-flattenable.
1.2.1. Example
@@
Figure 6. Graphs with one nonedge shown as a dashed-line segment. The graph-nonedge pair on the left has the $2$-SIP, while the one on the right does not. This illustrates that $d$-convexity is not closed under edge deletions. See below and also Section 2.
@@
We examine applications of Theorems 1 - 3 to the graph-nonedge pairs $(G,f)$ (left) and $(G^{\prime},f)$ (right) in Figure 6. The only atom of $G\cup f$ that contains $f$ is $K_{3}$, which does not have a $K_{4}$ minor. On the other hand, $G^{\prime}\cup f$ is an atom with a $K_{4}$ minor. Hence, Theorem 2 shows that $(G,f)$ has the $2$-SIP while $(G^{\prime},f)$ does not. This theorem also shows that neither of these pairs has the $1$-SIP. These statements can be verified using the techniques in Example 1.0.1.
@@
Next, $G\cup f$ is not $2$-flattenable since it has a $K_{4}$ minor. Therefore, Theorem 3 says that there exists some subset $F\subseteq E(G\cup f)$ such that $((G\cup f)\setminus F,F)$ is not $2$-convex. To verify this, let $f^{\prime}$ be the edge of $G$ deleted to obtain $G^{\prime}$. Since some atom of $G\cup f$ contains $f^{\prime}$ and is $K_{4}$, Theorem 2 shows that $((G\cup f)\setminus f^{\prime},f^{\prime})$ does not have the $2$-SIP.
The following provide motivation to study these properties. If a graph-nonedge-set pair $(G,F)$ is $d$-convex, then the CCS $\Omega^{d}_{F}(G,\ell)$ can be efficiently traversed and sampled using well-known methods. Additionally, if $G$ belongs to certain large and natural classes of graphs, then there are efficiently computable maps from $\Omega^{d}_{F}(G,\ell)$ to $\mathcal{C}^{d}(G,\ell)$ [24]. Furthermore, characterizations of $d$-convexity have led to improvements in molecular and particle assembly modeling and kinematic mechanism analysis and design, respectively, via the opensource software EASAL [15, 16] and CayMos [24, 28], and to efficient algorithms for the core problem of distance constraint graph realization [2].

### v2

# Graphs with single interval Cayley configuration spaces in 3-dimensions11footnotemark: 1

@@
Abstract
We prove a conjectured graph theoretic characterization of a geometric property of 3 dimensional linkages posed 15 years ago by Sitharam and Gao, motivated by their equivalent characterization for $d\leq 2$ that does not generalize to $d\geq 3$. A linkage $(G,\ell)$ contains a finite simple undirected graph $G$ and a map $\ell$ that assigns squared Euclidean lengths to the edges of $G$. A $d$-realization of $(G,\ell)$ is an assignment of points in $\mathbb{R}^{d}$ to the vertices of $G$ for which pairwise squared distances between points agree with $\ell$. For any positive integer $d\leq 3$, we characterize pairs $(G,f)$, where $f$ is a nonedge of $G$, such that, for any linkage $(G,\ell)$, the lengths attained by $f$ form a single interval - over the (typically a disconnected set of) $d$-realizations of $(G,\ell)$.
Although related to the minor closed class of $d$-flattenable graphs, the class of pairs $(G,f)$ with the above property is not closed under edge deletions, has no obvious well quasi-ordering, and there are infinitely many minimal graph-nonedge pairs - with respect to edge contractions - in the complement class. Our characterization overcomes these obstacles, is based on the forbidden minors for $d$-flattenability for $d\leq 3$, and contributes to the theory of Cayley configurations with many applications. Helper results and corollaries provide new tools for reasoning about configuration spaces and completions of partial 3-tree linkages, (non)convexity of Euclidean measurement sets in $3$-dimensions, their projections, fibers and sections. Generalizations to higher dimensions and efficient algorithmic characterizations are conjectured.
keywords:
@@ Abstract
Graph minors , Graph flattenability , Forbidden minor characterization , Discrete geometry , Distance geometry , Convex geometry , Configuration spaces , Cayley configuration spaces ,
1   Introduction
@@
A linkage $(G,\ell)$ is a pair containing a (finite simple undirected) graph $G$ and a squared edge-length map $\ell$ that sends each edge in the edge set $E(G)$ to a nonnegative real number. For any dimension $d\geq 0$, a (Euclidean) $d$-realization $p$ of $(G,\ell)$ is a map that sends each vertex in the vertex set $V(G)$ to a point in $\mathbb{R}^{d}$ such that $\ell(uv)=\|p(u)-p(v)\|^{2}$ for each edge $uv\in E(G)$. We omit “$d$” from this terminology when it is clear from context. The configuration space (CS) $\mathcal{C}^{d}(G,\ell)$ is the set of all $d$-realizations of $(G,\ell)$. Characterizing and sampling $\mathcal{C}^{d}(G,\ell)$ are difficult and well-studied problems [1, 2, 3, 4, 5, 6, 7] with applications in many domains [8, 9, 10].
Figure 1: Graph-nonedge pairs. See Example 1 and the discussion in this section, Example 3 and the discussion in Section 2, and the discussions in Sections 2.1 and 5.
@@
Figure 2: Left: linkages $(G,\ell)$ and sets $F$ of nonedges (dashed line-segments). Top-right: the CCS $\Omega^{3}_{F}(G,\ell)$ for the top-left linkage. Top-right recreated from [11]. Bottom-right: the CCS $\Omega^{2}_{F}(G,\sqrt{\ell})$ for the bottom-left linkage. Bottom recreated from [12]. See Example 1 and the discussion in this section, as well as Example 3 in Section 2.
@@
To tackle these problems, the paper [12] began characterizing graphs $G$ for which, for any $\ell$, the CS $\mathcal{C}^{d}(G,\ell)$, while typically disconnected, has a (branched) covering map to a base space that is convex. The notable feature is a graph theoretic characterization of a geometric property of linkages. For such graphs, the difficulty of sampling the CS in applications reduces to a much easier problem of sampling a convex base “configuration” space and inverting the covering map. More precisely, let $F=\{u_{1}v_{1},\dots,u_{m}v_{m}\}$ be an indexed set of nonedges of $G$, i.e., pairs of distinct vertices not in $E(G)$. The Cayley map $\phi_{F}$ sends a realization $p$ in $\mathcal{C}^{d}(G,\ell)$ to the vector $(\|p(u_{1})-p(v_{1})\|^{2},\dots,\|p(u_{m})-p(v_{m})\|^{2})$ of squared lengths attained by $F$ under $p$. The Cayley configuration space (CCS) $\Omega^{d}_{F}(G,\ell)$ is the image of $\mathcal{C}^{d}(G,\ell)$ under $\phi_{F}$. Note that the preimage of any point in $\Omega^{d}_{F}(G,\ell)$ under $\phi_{F}$ is a non-empty subset of $\mathcal{C}^{d}(G,\ell)$. See Figure 2 for examples. The pair $(G,F)$ is $d$-convex (resp. $d$-Cayley-connected) if $\Omega^{d}_{F}(G,\ell)$ is convex (resp. connected) for any squared edge-length map $\ell$.
Note 1.
In this paper, all sets are closed and compact, all maps are closed, and “connected” means “path-connected” unless otherwise specified. Also, $(G,F)$ is trivially $d$-convex and $d$-Cayley-connected if $d=0$, hence we assume that $d$ is an integer greater than 00, unless stated otherwise.
Characterizations of $d$-convexity for special classes of graphs have led to improvements in molecular and particle assembly modeling and kinematic mechanism analysis and design, respectively, via the opensource software EASAL [13, 8] and CayMos [14, 15], and to efficient algorithms for distance constraint graph realization [16].
We study the following property as a first step towards completely characterizing $d$-convex pairs $(G,F)$. If $F$ contains exactly one nonedge $f$, then we call $(G,f)$ a graph-nonedge pair.
@@
For any dimension $d\geq 1$, a graph-nonedge pair has the $d$-single interval property (SIP) if it is $d$-convex (equivalently, $d$-Cayley-connected).
The main result of this paper is a characterization for any $d\leq 3$ of graph-nonedge pairs $(G,f)$ with $d$-SIP, proving a conjecture posed in [12], which proved an equivalent characterization for $d\leq 2$ that does not generalize to $d\geq 3$.
Example 1.
Refer to the graph-nonedge pairs $(G,f)$ in Figure 1. If $G$ is the graph in Figure 1(a) without the blue edges, then there exists an $\ell$ for which both the CS $\mathcal{C}^{1}(G,\ell)$ (modulo translations) and CCS $\Omega^{1}_{f}(G,\ell)$ are finite non-singleton sets and hence $(G,f)$ does not have 1-SIP. On the other hand, the CS $\mathcal{C}^{2}(G,\ell)$ and CCS $\Omega^{2}_{f}(G,\ell)$ are clearly connected and hence $(G,f)$ has 2-SIP. Adding in the blue edges with almost any lengths causes $\mathcal{C}^{2}(G,\ell)$ to become disconnected while, for any lengths, all the connected components of $\mathcal{C}^{2}(G,\ell)$ induce the same interval $\Omega^{2}_{f}(G,\ell)$, and hence $(G,f)$ has 2-SIP.
If $G$ is the graph in Figure 1(b) without the blue edges, then there exists an $\ell$ for which both the CS $\mathcal{C}^{2}(G,\ell)$ (modulo translations and rotations) and CCS $\Omega^{2}_{f}(G,\ell)$ are finite non-singleton sets and hence $(G,f)$ does not have 2-SIP. On the other hand, the CS $\mathcal{C}^{3}(G,\ell)$ and CCS $\Omega^{3}_{f}(G,\ell)$ are clearly connected and hence $(G,f)$ has 3-SIP. Adding in the blue edges with almost any lengths causes $\mathcal{C}^{3}(G,\ell)$ to become disconnected (modulo translations and rotations) while, for any lengths, all the connected components of $\mathcal{C}^{3}(G,\ell)$ induce the same interval $\Omega^{3}_{f}(G,\ell)$, and hence $(G,f)$ has 3-SIP.
For similar reasons, if $G$ is the graph in Figure 1(d), $(G,f)$ has the $d$-SIP for any $d\geq 4$ but not for any $d<4$.
The linkage in Figure 1(c) also has the property that $\mathcal{C}^{3}(G,\ell)$ is disconnected. If we fix one of the two reflections of each of the 3 tetrahedra, $3$-SIP ($3$-convexity) of $(G,f^{\prime})$ is intuitively clear when restricted to this choice of reflections because the configuration space is connected. Although the different reflections could result in disconnected components of the CS $\mathcal{C}^{3}(G,\ell)$, with some effort it can be shown that all these components have the same attainable lengths for the nonedge $f^{\prime}$, which is sufficient (although not necessary) to prove $\Omega^{3}_{f^{\prime}}(G,\ell)$ is connected and hence $(G,f^{\prime})$ has 3-SIP. Showing a weaker version of this property in general is sufficient to show $3$-SIP for $(G,f)$ e.g. in Figure 1(c) (see also Figure 12(d)). This is a key property established in the converse direction of Theorem 4 in Section 2 - our main result.
@@
Next, refer to the pairs $(G,F)$ in Figure 2. For the linkage $(G,\ell)$ in the bottom-left image, the CCS $\Omega^{2}_{F}(G,\sqrt{\ell})$ shown in the bottom-right image is the intersection of half-spaces determined by triangle-inequalities on $\sqrt{\ell}$, and it is not hard to see that $\Omega^{2}_{F}(G,\ell)$ is also convex. Since this is true for any $\ell$, $(G,F)$ is $d$-convex for any $d\geq 2$. On the other hand, by choosing $\ell$ so that one edge has length 00 and the rest have length $1$, we get that $\Omega^{1}_{f}(G,\ell)$ has cardinality two, and so $(G,F)$ is not $1$-convex. In the top-left image, a similar argument using tetrahedral-inequalities shows that $(G,F)$ is $d$-convex for any $d\geq 3$ but not for any $d<3$. The top-right image shows the convex CCS $\Omega^{3}_{F}(G,\ell)$ for the linkage $(G,\ell)$ in the top-left image. The CS $\mathcal{C}^{3}(G,\ell)$ is clearly connected for any $\ell$, and so $\Omega^{3}_{F}(G,\ell)$ is connected for any $\ell$ since $\phi_{F}$ is continuous. Therefore, $(G,F)$ is $2$-Cayley-connected. However, an argument similar to those above shows that $(G,F)$ is not $1$-Cayley-connected.
@@
Figure 3: The graphs $C_{5}\times C_{2}$ (left) and $V_{8}$ (right) along with nonedge sets (shown as dashed line-segments) that together are $d$-convex pairs for any dimension $d\geq 3$ but not for any $d<3$. See Example 1 in this section, Example 3 and the discussion in Section 2, and Example 6 in Section 5.
@@
Finally, the pairs $(G,F)$ in Figure 3 are $d$-convex for any dimension $d\geq 3$ but not for any $d<3$. The case $d\geq 3$ can be verified for the first two pairs using Theorem 3 in Section 1.2. The case $d<3$ can be verified intuitively, and using Theorems 1 and 2 in Section 1.2. ∎
1.1   Organization
@@
Previous results on $d$-conxevity, $d$-Cayley-connectedness, and the $d$-SIP are collected in Section 1.2. We state our contributions in Section 2, including Theorem 4, our main result, which gives a characterization of graph-nonedge pairs that have the $3$-SIP; and several frequently used and independently interesting tool lemmas, byproducts, and corollaries, which are proved in Sections 4.1 and 4.2. The majority of this paper is spent proving Theorem 4. The primary proof obstacles are detailed in Section 2.1. Since our proof techniques involve very technical work on graph minors, notation and terminology to facilitate this is introduced in Section 3. The proof is given in Sections 4 (converse direction) and 5 (forward direction). Finally, we state several open problems in Section 6, including a conjectured characterization of graph-nonedge pairs that have the $d$-SIP for any $d\geq 1$. The appendix collects proofs of independently interesting results about clique separators, partial 3-trees, and related covering maps, and long and tedious but not particularly illuminating proofs for helper lemmas.
1.2   Related Works
@@
Theorems 1 and 2, below, are complete characterizations of $d$-convexity and the $d$-SIP for any $d\leq 2$, given in [12] and restated using our terminology. The proof of the former strongly relies on the latter. See Example 1, in Section 1, and Example 2, below, for applications of these theorems. We require the following concepts to state these theorems. Given a graph $G$ and a set of its edges and nonedges $F$, $G\cup F$ is the graph obtained from $G$ by adding $F$ to $E(G)$. A subgraph $H$ of $G$ is induced if, for any $u,v\in V(H)$, $uv$ is an edge of $H$ if and only if it is an edge of $G$. Furthermore, $H$ is induced by a set $U\subseteq V(G)$ if $H$ is induced and $V(H)=U$. The subgraph induced by $V(G)\setminus U$ is written as $G\setminus U$. If either $F$ or $U$ contains exactly one element $f$ or $u$, respectively, then we replace $F$ with $f$ and $U$ with $u$ in this notation. $U$ is a separator of $G$ if $G$ is connected and $G\setminus U$ is disconnected, a clique of $G$ if its induces a clique subgraph, and a clique separator of $G$ if it is both a clique and a separator. An atom is a graph with no clique separator. An atom of $G$ is a vertex-maximal induced subgraph with no clique separator. See Figure 4 for examples.
Figure 4: (a) a graph-nonedge pair $(G,f)$ with $G\cup f$ having a clique separator $C$, shown in purple, (b) the $C$-components of $G\cup f$, and (c) the atoms of $G\cup f$. See the discussion in this section, Example 3 in Section 2, and Example 6 in Section 5.
A graph $M$ is a minor of $G$ if it is isomorphic to some graph obtained from $G$ via some sequence of edge deletions and (edge) contractions - i.e., replacing two adjacent vertices by a single vertex whose neighborhood is the union of their neighborhoods.
@@
For any dimension $d\leq 2$, and graph $G$ with nonedge set $F$, the pair $(G,F)$ is $d$-convex if and only if each atom of $G\cup F$ that contains a non-empty subset of $F$ has no $K_{d+2}$ minor.
@@
For any dimension $d\leq 2$, a graph-nonedge pair $(G,f)$ has the $d$-SIP if and only if each atom of $G\cup f$ that contains $f$ has no $K_{d+2}$ minor.
The above theorem does not directly generalize to $d\geq 3$ as discussed in more detail in Example 3 in Section 2 (see Figure 8 for a counterexample). Hence the following modification was conjectured in [12] for $d=3$, 15 years ago, and is the main result in this paper.
Conjecture 1 ([12], Conjecture 6.4).
A graph-nonedge pair $(G,f)$ has the $3$-SIP if and only if $f$ must be contracted in any $K_{5}$ or $K_{2,2,2}$ minor of any atom of $G\cup f$ that contains $f$.
The above minors are central to another property of graphs called $d$-flattenability, introduced in [17, 18]. A graph $G$ is $d$-flattenable if, for any squared edge-length map $\ell$ and any dimension $n\geq d$, the CS $\mathcal{C}^{d}(G,\ell)$ is non-empty whenever the CS $\mathcal{C}^{n}(G,\ell)$ is non-empty. Intuitively, this means that any high-dimensional realization of any linkage $(G,\ell)$ can be flattened to live in $d$-dimensions while preserving squared edge-lengths. For example, $K_{3}$ is clearly $2$-flattenable but not $1$-flattenable, since choosing a squared edge-length map $\ell$ that assigns $1$ to each edge of $K_{3}$ ensures the CS $\mathcal{C}^{1}(K_{3},\ell)$ is empty even though the CS $\mathcal{C}^{2}(K_{3},\ell)$ is not.
Figure 5: The $3$-flattenability forbidden minors: $K_{5}$ (left) and $K_{2,2,2}$ (right). See the discussions in this section and Section 2.
A property of graphs is minor closed if it is preserved under edge deletions and contractions. More precisely: given a graph $G$ with the property, any graph obtained by deleting or contracting edges of $G$ also has the property. In other words, the set of graphs that have the property is closed under edge deletions and contractions. It is easy to see that $d$-flattenability is minor closed. Hence, Robertson and Seymour’s famous Graph Minor Theorem [19] guarantees the existence of a finite set of forbidden minors such that a graph is $d$-flattenable if and only if it has no minor in this set. It is folklore that $K_{d+2}$ is the only graph in this set for $d\in\{1,2\}$, and proofs can be found in [17], where they proved that $K_{5}$ and $K_{2,2,2}$ are the only graphs in this set for $d=3$ (see Figure 5). Finite forbidden minor characterizations for $d$-flattenability have also been studied for normed spaces other than Euclidean [20, 21, 22]. With these characterizations, the right-hand side of Theorem 2 is equivalent to “each atom of $G\cup f$ that contains $f$ has no $d$-flattenability forbidden minor”. As mentioned earlier, we will see in Example 3, in Section 2, that this restatement does not extend to $d=3$ (see Figure 8 for a counterexample).
A relationship was established in [12] between $d$-convexity and $d$-flattenability for $d=3$, and later extended in [23] to any $d\geq 0$ and to other norms. This relationship makes use of previous work on understanding the Euclidean and non-Euclidean distance cones [24, 25]. Given a graph $G$ and a set $F$ of its edges and nonedges, $G\setminus F$ is the graph obtained from $G$ by deleting $F$ from $E(G)$. If $F$ contains exactly one element $f$, then we replace $F$ with $f$ in this notation.
@@
For any dimension $d\geq 0$ and any graph $G$ the following are equivalent:
  1. 1.
    $G$ is $d$-flattenable.
  2. 2.
    for any subset $F\subseteq E(G)$, the pair $(G\setminus F,F)$ is $d$-convex.
@@
Figure 6: The graph-nonedge pair on the left has the $2$-SIP, while the one on the right does not (nonedges are shown as dashed-line segments). This illustrates that $d$-convexity is not closed under edge deletions. See Example 2 in this section, Examples 3 and 5 in Section 2, and the discussion in Section 2.1.
Example 2.
@@
We examine applications of Theorems 1 - 3 to the graph-nonedge pairs $(G,f)$ (left) and $(G^{\prime},f)$ (right) in Figure 6. The only atom of $G\cup f$ that contains $f$ is $K_{3}$, which does not have a $K_{4}$ minor. On the other hand, $G^{\prime}\cup f$ is an atom with a $K_{4}$ minor. Hence, Theorem 2 shows that $(G,f)$ has the $2$-SIP while $(G^{\prime},f)$ does not. This theorem also shows that neither of these pairs has the $1$-SIP. These statements can be verified using the observations in Example 1.
@@
Next, $G\cup f$ is not $2$-flattenable since it has a $K_{4}$ minor. Therefore, Theorem 3 says that there exists some subset $F\subseteq E(G\cup f)$ such that $((G\cup f)\setminus F,F)$ is not $2$-convex. To verify this, let $f^{\prime}$ be the edge of $G$ deleted to obtain $G^{\prime}$. Since some atom of $G\cup f$ contains $f^{\prime}$ and is $K_{4}$, Theorem 2 shows that $((G\cup f)\setminus f^{\prime},f^{\prime})$ does not have the $2$-SIP. ∎